the speed of a stream is 2 mph. a boat travels 4 miles upstream in the same time it take to travel 8 miles downstream. what is the speed of the boat in still water?
Answer by amarjeeth123(569)

the speed of a stream is 2 mph. a boat travels 4 miles upstream in the same time it take to travel 8 miles downstream.
Time=Distance/Speed
Let the speed of the boat in still water be x.
Time taken to travel upstream=4/(x-2)
Time taken to travel downstream=8/(2+x)
4/(x-2)=8/(2+x)
Simplifying we get,
2+x=2x-4
x=6 miles/hour
Speed of the boat in still water=6 miles/hour.